1.) In this experiment, if the carboxylic acid is benzoic acid, how many moles of benzoic acid are present (assume an equal portion of each component)? How many moles of sodium bicarbonate are contained in 1 mL of a saturated aqueous sodium bicarbonate. Is the amount of sodium bicarbonate sufficient to react with all of the benzoic acid?

molar mass of Benzoic acid =122 gm, moles of benzoic acid in 0.1385 gm= mass/molar mass= 0.1385/122= 0.001135
Answer #1

1 mole of sodium bicarbonate reacts with 1 mol of benzoic acid.

Solubility of NaHCO3 = 69 g / Liter = 0.069 g / ml

Moles of NaHCO3 = mass / molar mass = 0.069 / 84 = 8.21 * 10-4 mol

This si not enough to react with all of the benzoic acid